OpenStudy (linn99123):

If h(z)= (0.5)(z-7)(z+7), find h(-3)

OpenStudy (phi):

h(z)= (0.5)(z-7)(z+7) the first step to find h(-3) is erase the "z" and put in -3 can you do that? what do you get ?

OpenStudy (phi):

you should get h(-3)= 0.5(-3-7)(-3+7) remember order of operations: do the stuff in parens first: -3-7= -10 and -3+7= 4 so h(-3) = 0.5*-10*4 now multiply the 3 numbers.